OpenStudy (anonymous):

Solve this problem on paper using all four steps. Find four consecutive integers whose sum is 114. The integers are , , , and .

OpenStudy (anonymous):

do 114/4=28.5 then pick 2 numbers greater than and 2 less than in a row 27+28+29+30=114

OpenStudy (haseeb96):

let suppose n is the interger four consecutive integers are n, n+1,n+2,n+3 solution :- n+ n+1+n+2+n+3=114 4n +6=114 4n =114-6 4n = 108 n = 108 /4 n= 27 so four consecutive integers will be 27, 28, 29, 30 Answer

OpenStudy (anonymous):

@jjmack2 dont forget to medal and fan those who help u

OpenStudy (michele_laino):

hint: your problem can be modeled by the subsequent equation: \[\Large n + \left( {n + 1} \right) + \left( {n + 2} \right) + \left( {n + 3} \right) = 114\]

OpenStudy (michele_laino):

where n is an integer number

OpenStudy (michele_laino):

please, solve that equation for n